# Distributed and heterogeneous tensor–vector contraction algorithms for high performance computing

### v1

@@ 4.2 Distributed-memory higher order power method
In the next paragraphs we analyse the HOPM in terms of intra-process data movement exclusively. In the particular case of hypersquare tensors ($n_{i}=n,$), a sequential implementation of the HOPM that can be found in Refs. [3, 6] incurs a certain amount of touched memory
@@
In contrast to Eq. (3), the touched memory expression of the distributed HOPM algorithm (dHOPM) loses the symmetry of its sequential counterpart. Each parallel process incurs $M_{\rm par}=sm_{{\rm par},j<s}+m_{{\rm par},j=s}+(d-1-s)m_{{\rm par},j>s}$, with
@@
in the particular case of $j=s$. Square brackets denote the optimal division but, for the sake of simplicity, we assume a regular division thereby approximating the previous equation so that it can be related to Eq. (3). In fact, the second term of Eq. (4.2) reveals that the duplication of the input vectors on each MPI process causes a memory overhead over the sequential algorithm. For the remaining $d-1$ external iterations, we get the following expression
@@
with $l=0$ if $j<s$, otherwise $l=1$ if $j>s$. It can be noticed that Eq. (4.2) incurs further overhead over Eq. (4.2) because, contrary to the previous case, now the contraction mode and the splitting index will coincide ($k=s$) with one of the $d-1$ inner TVCs thus yielding larger partial subtensors as discussed in Section 4.1. This is indeed a very interesting aspect of the dHOPM algorithm: it will lead to $d-1$ underperforming contractions. The final expression for the dHOPM touched memory per process can be written as
@@
that elucidates a linear increase of the data movement with the number of processes and a more complex, nonlinear relationship with the splitting dimension. To summarize, one should avoid splitting tensors along their first dimensions, especially when using many processes. This will reduce the memory streamed by the distributed HOPM and, due to the memory-bound nature of this particular algorithm, it will translate into faster computations.
@@
Let us consider the following nondimensional variables: $\hat{p}=p/n$, $\hat{s}=s/(d-1)$, and $\eta^{-1}(\hat{p},\hat{s})=pM_{\rm par}/M_{\rm seq}$. Figure 2(a) shows the memory ratio $\eta^{-1}$ of the classical dHOPM for the third- and tenth-order hypersquare tensors considered in this work (see Table 1). In both cases, the data movement more than doubles for $\hat{s}=0$ and $\hat{p}=1$. With $d=3$ the memory ratio increases steadily for decreasing values of $\hat{s}$ while, for $d=10$, memory effects are limited to a narrow region of low values of $\hat{s}$. As expected, the touched memory tends to augment with the number of processes and, in any case, if the tensor is split along its last dimension, then $M_{\rm par}\approx M_{\rm par,min}\approx M_{\rm seq}/p$. A similar trend is observed for the other tensors of Table 1 which are not shown in Fig. 2(a) for the sake of conciseness. The only exception to this rule comes from the second-order tensor (matrix) where $M_{\rm par,min}$ only varies with $p$, as it can be deduced from Eq. (6), and its value can be significantly larger than those associated with higher order tensors. Finally, it is worth noting that the expressions (3)–(7) and Fig. 2 are still representative of nonhypersquare tensors provided that the size assigned to each dimension remains about the same order of magnitude.
@@
3             $\alpha\leftarrow 0$ if $j>0$, else $\alpha\leftarrow 1$ ;
4             $\beta\leftarrow\max(\alpha,\,j)$ ;
5             $\gamma\leftarrow d-1$ if $j<d-1$, else $\gamma\leftarrow d-2$ ;
@@
7                   $\mathcal{W}_{j}^{(p)}:=\mathcal{Y}_{\alpha}^{(p)}\;\leftarrow\mathcal{A}^{(p)}\times_{\alpha}\mathbf{x}_{\alpha}$ ;
@@
10                   $\mathcal{W}_{j}^{(p)}:=\mathcal{Y}_{\beta-1}^{(p)}\;\leftarrow\mathcal{W}_{j-1}^{(p)}\times_{\beta-1}\mathbf{x}_{\beta-1}$ ;
@@
            $\mathcal{Y}_{\beta+1}^{(p)}\leftarrow\mathcal{Y}_{\beta+\alpha-1}^{(p)}\times_{\beta+1}\mathbf{x}_{\beta+1}$ ;
@@ 13                   $\mathcal{Y}_{k}^{(p)}\leftarrow\mathcal{Y}_{k-1}^{(p)}\times_{k}\mathbf{x}_{k}$ ;
                   $\mathbf{x}_{j}\leftarrow\sum_{p}\mathcal{Y}_{\gamma}^{(p)}$ ;
@@
                   $\mathbf{x}_{j}\leftarrow\bigsqcup_{p}\mathcal{Y}_{\gamma}^{(p)}$ ;
@@
Algorithm 1 Distributed-memory, three-buffer higher order power method algorithm (dHOPM3).
@@
The classical HOPM algorithm can be further optimized for distributed systems as shown in Algorithm 1. This novel approach, which makes use of three buffers, hence dHOPM3, uses two buffers $\mathcal{Y}_{k-1}$ and $\mathcal{Y}_{k}$ to save intermediate tensors and a third buffer $\mathcal{W}$ to hold previously computed data in memory ($k=\alpha$ or $k=\beta-1$, see lines 7 and 9) for the next iteration when $j\geq 2$ (line 9). This allows us to save a total of $(d-1)(d-2)/2$ contractions and, as the order of the original tensor keeps increasing, dHOPM3 can save up to half of the total contractions required by a canonical, two-buffer algorithm such as the one from Pawłowski et al. [6]. What is more important, the skipped contractions are precisely the most computationally expensive ones since they are carried out on the largest input tensors, which results in even better computational speedups. Another aspect of Algorithm 1 worth mentioning is that it inherently exploits the distributive property of consecutive TVCs and, consequently, the summation required by Eq. (2) when $j\neq s$ is delayed until the normalization step, thereby reducing the communication to only $n_{j}$ elements per process at line 14; in other words, the consecutive TVC operations at lines 6–12 are done exclusively on partial, distributed subtensors without incurring in global assembly penalizations. Alternatively, the case $j=s$ implies the disjoint union of $\mathbf{x}_{j}$ (line 16) where each process exchanges their calculated portion $\mathbf{x}_{j}^{(p)}$ for an aggregated communication of $n_{j}$ elements per process. Finally, the resulting vector is normalized locally by each process at line 17.
@@
Now we can define $H^{-1}(\hat{p},\hat{s})$ as the streamed memory ratio between the classical distributed HOPM and our optimal distributed implementation dHOPM3. Figure 2(b) shows the evolution of $H^{-1}$ for the previously analyzed third- and tenth-order tensors. We obtain two approximately flat surfaces, especially for the low order tensor, which indicates that this ratio is almost independent of $\hat{p}$ and $\hat{s}$. Our optimized algorithm economizes about 1.5$\times$ of the touched memory for $d=3$ and roughly a fivefold for $d=10$ (with the presence of a minimum of about 3.3$\times$). All things considered, Figs. 2(a)–(b) demonstrate that dHOPM3 can save up to one order of magnitude of streamed memory for higher order tensors in comparison to a canonical HOPM implementation with naively split tensors.
@@
As previously mentioned, Eqs. (4.2)–(7) do not account for inter-process data movement. On the one hand, for all the external iterations except one, the resulting vector $\mathbf{x}_{j}$ must be entirely reduced by all processes (see line 14 of Algorithm 1), otherwise only a portion of this array is gathered (line 16). The actual amount of data being both transferred and computed ultimately depends on the algorithm employed underneath the functions MPI_Allgather() and MPI_Allreduce() and it is likely to change at runtime based on the message size, the number of processes and the network topology [27]. For example, the straightforward Ring algorithm, which is bandwidth-optimal, suitable for large messages, and works with any value of $p$, yields $4n(p-1)/p$ extra touched memory per process, a quantity that can be appended to the term $M_{\rm par,min}$ in Eq. (6). It can be easily inferred that the inter-process contribution reaches its maximum for $d=2$ and $\hat{p}=p/n=1$, increasing the value of $M_{\rm par,min}$ by up to $4/7\approx 57\%$. In the case of hypersquare tensors, this constant number remains independent of the splitting dimension and can be halved by storing separate copies of the input tensor $\mathcal{A}$ with different splittings in order to foster partial gatherings of $\mathbf{x}_{j}$.
@@ 5 Performance evaluation
This section evaluates the performance of the distributed TVC and HOPM algorithms detailed in previous sections of this manuscript after describing the numerical environment setup. The complete source code of this distributed TVC library will be made publicly available under the GPLv3 license upon the acceptance of this manuscript [28, 29].
@@ 5.1 HPC systems and code setup
The tensors used in this work are listed in Table 1. We choose hypersquare and up to tenth-order dense tensors. In principle, they are filled with 8-byte floating-point numbers (i.e., doubles) following other works [6, 8]. Lower precision floats will be introduced later in Section 5.5. When possible, each tensor dimension size has been intentionally chosen to prevent it from being a multiple of the vector length —which is 8 when using doubles and the 512-bit SIMD/SVE instruction set— with the aim of fostering scenarios showcasing unaligned memory accesses as well as peeling and remainder loops. The aforementioned tensors average 7.5 GB of memory in order to guarantee that TVC operations are limited by the system memory bandwidth while fitting within the rather small HBM2 memory incorporated in two of the three systems described above. Consequently, larger tensors with vector-friendly dimension sizes are expected to yield better performance figures.
With regard to code compilation, we utilize the LLVM infrastructure unless otherwise stated. In particular, we employ the Clang++ 16.0 compiler that supports both the OpenMP [30] and OmpSs-2 [31] programming languages allowing for data-flow parallelization via tasks. In order to obtain the maximum performance, we use the flag -march= native to enable CPU specific optimizations. The combination of flags -Ofast and -mprefer-vector-width=512 are used to exploit 512-bit vector instructions. In addition to this, the flag -fopenmp enables OpenMP pragma directives used for both vectorization and parallelization of the code: unless otherwise stated, shared-memory parallelization is achieved via fork-join, “parallel for” loops. Furthermore, the flag -flto enables interprocedural (i.e., link-time) optimizations. Other code parameters are the alignment of buffers to transparent huge pages of 2 MiB and the loop unrolling factor set to 8. For GPU devices, we make use of the latest cuBLAS 10.2 library officially supported by the architecture vendor.

### v2

@@ 4.2 Distributed-memory higher-order power method
In the next paragraphs we analyse the HOPM in terms of intra-process data movement exclusively. In the particular case of hypersquare tensors ($n_{i}=n$), a sequential implementation of the HOPM that can be found in Refs. [3, 6] incurs a certain amount of touched memory
@@
In contrast to Eq. (3), the touched memory expression of the distributed HOPM algorithm loses the symmetry of its sequential counterpart. Each parallel process incurs $M_{\rm par}=sm_{{\rm par},j<s}+m_{{\rm par},j=s}+(d-1-s)m_{{\rm par},j>s}$, with
@@
when the external iteration $j$ coincides with the splitting dimension $s$. Square brackets denote the optimal division but, for the sake of simplicity, we assume a regular division thereby approximating the previous equation so that it can be related to Eq. (3). In fact, the second term of Eq. (4.2) reveals that the duplication of the input vectors on each MPI process causes a memory overhead over the sequential algorithm. For the remaining $d-1$ external iterations, we get the following expression
@@
with $l=0$ if $j<s$, otherwise $l=1$ if $j>s$. It can be noticed that Eq. (4.2) incurs further overhead over Eq. (4.2) because, contrary to the previous case, now the contraction mode and the splitting index will coincide ($k=s$) during one of the $d-1$ inner TVCs thus yielding larger partial subtensors as discussed in Section 4.1. This is indeed a very interesting aspect of the dHOPM algorithm: it will lead to $d-1$ underperforming contractions. The final expression for the dHOPM touched memory per process can be written as
@@
that elucidates a linear increase of data movement with the number of processes and a more complex, nonlinear relationship with the splitting dimension. Contrary to the dTVC algorithm and its later assembly process described at the end of Section 4.1, for the dHOPM one should avoid splitting tensors along their first dimensions, especially when using many processes. This will significantly reduce the movement of data and yield faster computations.
@@
Let us consider the following nondimensional variables: $\hat{p}=p/n$, $\hat{s}=s/(d-1)$, and $\eta^{-1}(\hat{p},\hat{s})=pM_{\rm par}/M_{\rm seq}$. Figure 2(a) shows the memory ratio $\eta^{-1}$ of the classical dHOPM for the third- and tenth-order hypersquare tensors considered in this work (see Table 1). In both cases, the data movement more than doubles for $\hat{s}=0$ and $\hat{p}=1$. With $d=3$ the memory ratio increases steadily for decreasing values of $\hat{s}$ while, for $d=10$, memory effects are limited to a narrow region of low values of $\hat{s}$. As expected, the touched memory tends to augment with the number of processes and, in any case, if the tensor is split along its last dimension, then $M_{\rm par}\approx M_{\rm par,min}\approx M_{\rm seq}/p$. A similar trend is observed for the other tensors of Table 1 which are not shown in Fig. 2(a) for the sake of conciseness. The only exception to this rule comes from the second-order tensor (matrix) where $M_{\rm par,min}$ only varies with $p$, as it can be deduced from Eq. (6), and its value can be significantly larger than those associated with higher-order tensors. Finally, it is worth noting that the expressions (3)–(7) and Fig. 2 are still representative of nonhypersquare tensors provided that the size assigned to each dimension remains about the same order of magnitude.
@@
3             $\lambda\leftarrow 0$ if $j>0$, else $\lambda\leftarrow 1$ ;
4             $\mu\leftarrow\max(\lambda,j)$ ;
5             $\nu\leftarrow d-1$ if $j<d-1$, else $\nu\leftarrow d-2$ ;
@@
7                   $\mathcal{Y}_{\lambda}^{(p)}:=\mathcal{W}_{j}^{(p)}\leftarrow\mathcal{A}^{(p)}\times_{\lambda}\mathbf{x}_{\lambda}$ ;
@@
10                   $\mathcal{Y}_{\mu-1}^{(p)}:=\mathcal{W}_{j}^{(p)}\leftarrow\mathcal{W}_{j-1}^{(p)}\times_{\mu-1}\mathbf{x}_{\mu-1}$ ;
@@
            $\mathcal{Y}_{\mu+1}^{(p)}\leftarrow\mathcal{Y}_{\mu+\lambda-1}^{(p)}\times_{\mu+1}\mathbf{x}_{\mu+1}$ ;
@@ 13                   $\mathcal{Y}_{k}^{(p)}\leftarrow\mathcal{Y}_{k-1}^{(p)}\times_{k}\mathbf{x}_{k}$ ;
                   $\mathbf{x}_{j}\leftarrow\sum_{p}\mathcal{Y}_{\nu\phantom{k}}^{(p)}$ ;
@@
                   $\mathbf{x}_{j}\leftarrow\bigsqcup_{p}\mathcal{Y}_{\nu\phantom{k}}^{(p)}$ ;
@@
Algorithm 1 Distributed-memory, three-buffer higher-order power method algorithm (dHOPM3).
@@
The classical dHOPM analyzed above can be further optimized for distributed systems as shown in Algorithm 1. This novel approach, which makes use of three buffers (dHOPM3), employs two of them $\mathcal{Y}_{k-1}$ and $\mathcal{Y}_{k}$ to save intermediate tensors and a third one $\mathcal{W}$ to hold previously computed data in memory ($k=\lambda$ or $k=\mu-1$, see lines 7 or 9) for the next iteration when $j\geq 2$ (line 9). This allows us to save a total of $(d-1)(d-2)/2$ contractions and, as the order of the original input tensor keeps increasing, dHOPM3 can save up to half of the total contractions required by a canonical, two-buffer algorithm such as the one from Pawłowski et al. [6]. What is more important, the skipped contractions are precisely the most computationally expensive ones since they are carried out on the largest input tensors, which results in even better computational speedups. Another aspect of Algorithm 1 worth mentioning is that it inherently exploits the distributive property of consecutive TVCs and, consequently, the summation required by Eq. (2) when $k=s$ (which implies $j\neq s$) is delayed until the normalization step, thereby reducing the communication to only $n_{j}$ elements per process at line 14; in other words, the consecutive TVC operations at lines 6–12 are done exclusively on partial, distributed subtensors without incurring in global assembly penalizations. Alternatively, the case $j=s$ implies the disjoint union of $\mathbf{x}_{j}$ (line 16) where each process exchanges their calculated portion $\mathbf{x}_{j}^{(p)}$ for an aggregated communication of $n_{j}$ elements per process. Finally, the resulting vector is normalized locally by each process at line 17.
@@
Now we can define $H^{-1}(\hat{p},\hat{s})$ as the streamed memory ratio between the classical dHOPM and our optimal implementation dHOPM3. Figure 2(b) shows the evolution of $H^{-1}$ for the previously analyzed third- and tenth-order tensors. We obtain two approximately flat surfaces, especially for the low-order tensor, which indicates that this ratio is almost independent of $\hat{p}$ and $\hat{s}$. Our optimized algorithm economizes about 1.5$\times$ of the touched memory for $d=3$ and roughly a fivefold for $d=10$ (with the presence of a minimum of about 3.3$\times$). All things considered, Figs. 2(a)–(b) demonstrate that dHOPM3 can save up to one order of magnitude of streamed memory for higher-order tensors in comparison to a canonical HOPM implementation with naively split tensors.
@@
As previously mentioned, Eqs. (4.2)–(7) do not account for inter-process data movement. On the one hand, for all the external iterations except one, the resulting vector $\mathbf{x}_{j}$ must be entirely reduced by all processes (see line 14 of Algorithm 1), otherwise only a portion of this array is gathered (line 16). The actual amount of data being both transferred and computed ultimately depends on the algorithm employed underneath the functions MPI_Allgather and MPI_Allreduce and it is likely to change at runtime based on the message size, the number of processes and the network topology [29]. For example, the straightforward ring algorithm, which is bandwidth-optimal, suitable for large messages, and works with any value of $p$, yields $4n(p-1)/p$ extra touched memory per process, a quantity that can be appended to the term $M_{\rm par,min}$ in Eq. (6). It can be easily inferred that the inter-process contribution reaches its maximum for $d=2$ and $\hat{p}=p/n=1$, increasing the value of $M_{\rm par,min}$ by up to $4/7\approx 57\%$. In the case of hypersquare tensors, this constant number remains independent of the splitting dimension and can be halved by storing separate copies of the input tensor $\mathcal{A}$ with different splittings in order to foster partial gatherings of $\mathbf{x}_{j}$.
(a) Full dHOPM3 execution ($i=0$)
(b) Zoom-in on the last $k$ internal contractions of iteration $j=0$
Figure 3: Paraver traces from the hybrid execution of (a) the dHOPM3 algorithm and (b) a zoom-in view of the last $k$ internal contractions of iteration $j=0$ using the tenth-order tensor of Table 1 split along its last dimension. Traces obtained on MN4 using 4 MPI processes, 8 cores per process, and 32 OmpSs-2 tasks per process for illustration purposes.
4.3 Task-based parallelization of the dHOPM3 algorithm
We take the opportunity to describe herein the shared-memory parallel strategy adopted for dHOPM3 in which every contraction (lines 7, 9, 10, and 12 of Algorithm 1) takes advantage of task-based parallelization. In contrast to ubiquitous OpenMP parallel for loops, tasks permit to overlap computations corresponding to different TVC instances, thereby maximizing the parallel performance over fork-join. This key novel parallel design is exemplified below for a single TVC operation, defined by the getvc function [10], which is split in several tasks:
  if (trans == iZero) { // Matrix-vector (row-major)
    for (intT i = iZero; i < m; i += bsM) {
      #pragma omp task depend(in : x[iZero:     n]) \
                       depend(in : a[i*lda:bsMA*n]) \
                       depend(out: y[i    :bsMA  ])
      getvc(layout, trans, bsMA, n,
            alpha, a + i*lda, lda,
            x, incx, y, beta, y + i, incy);
  } else { // Vector-matrix (row-major)
    for (intT j = iZero; j < n; j += bsN) {
      #pragma omp task depend(in : x[iZero :m   ]) \
                       depend(in : a[startl:size]) \
                       depend(out: y[j     :bsNA])
      getvc(layout, trans, m, bsNA,
            alpha, a + startl, lda,
            x, incx, y, beta, y + j, incy);
Each TVC is assigned a certain number of tasks (i.e., granularity) inside a loop that can be executed concurrently. We establish task data dependencies via depend clauses for the input vector x and tensor a as well as the output tensor y to ensure the correct order of execution of different TVCs and normalization steps as indicated in Algorithm 1. The task granularity is such that there are more tasks than threads, which permits the overlapping of subsequent TVC operations. This is clearly seen in Fig. 3, which corresponds to a real execution carried out on MN4, instrumented with Extrae and visualized with Paraver [^4] Extrae and Paraver are available at: https://tools.bsc.es/ for a tenth-order tensor distributed among four processes. While Fig. 3(a) covers a full dHOPM3 execution ($i=0$), the figure below is a zoomed version towards the end of the first external iteration $j=0$. In both figures, subsequent TVC operations are represented with rectangles (tasks) of different colors. As previously mentioned, threads can start executing blocks of different colors as quickly as possible. There are still some relatively small white gaps that evidence code sections with no CPU usage and correspond exclusively to the final normalization phase (lines 13–16 of Algorithm 1) where synchronous collective communication takes place and MPI processes must wait for the slowest one due to system imbalance. Both subfigures also reveal how computational costs greatly reduce after each contraction as the resulting tensor loses one dimension and its size decreases by about one order of magnitude.
Contrary to tasks, a shared-memory strategy based upon the fork-join paradigm (see the code snippet from Ref. [8]) and, by extension a CUDA-based implementation, presents one implicit synchronization per contraction. Consequently, a Paraver trace of such parallel execution (not shown here for the sake of conciseness) would have shown white gaps towards the end of each TVC where idle threads must wait to the slowest one to finish its corresponding computation, close the parallelism (join) and, finally, reopen it for the next contraction (fork). All these additional periods of CPU inactivity increment the computational time and the application imbalance that ultimately result in an overall lower speedup with respect to the task-based implementation.
@@ 5 Performance evaluation
This section evaluates the performance of the distributed TVC and HOPM algorithms detailed in previous sections of this manuscript after describing the numerical environment setup. The complete source code of our dTVC library is made publicly available under the GPLv3 license in Ref. [10].
@@ 5.1 HPC systems and code setup
The tensors used in this work are listed in Table 1. We choose hypersquare and up to tenth-order dense tensors. In principle, they are filled with 8-byte floating-point numbers (i.e., doubles) following other works [6, 8]. Lower-precision floats will be introduced later in Section 5.5. When possible, each tensor dimension size has been intentionally chosen to prevent it from being a multiple of the vector length —which is 8 when using doubles and the 512-bit SIMD/SVE instruction set— with the aim of fostering scenarios showcasing unaligned memory accesses as well as peeling and remainder loops. The aforementioned tensors average 7.5 GB of memory in order to guarantee that TVC operations are limited by the system memory bandwidth while fitting within the rather small HBM2 memory incorporated in two of the three systems described above. Consequently, larger tensors with vector-friendly dimension sizes are expected to yield better performance figures.
With regard to code compilation, we utilize the LLVM infrastructure unless otherwise stated. In particular, we employ the Clang++ 16.0 compiler that supports both the OpenMP [30] and OmpSs-2 [31] programming languages allowing for data-flow parallelization via task annotations. In order to obtain the maximum performance, we use the flag -march= native to enable CPU specific optimizations. The combination of flags -Ofast and -mprefer-vector-width=512 are used to exploit 512-bit vector instructions. In addition to this, the flag -fopenmp enables OpenMP pragma directives used for both vectorization and parallelization of the code: unless otherwise stated, shared-memory parallelization is achieved via fork-join, “parallel for” loops. Furthermore, the flag -flto enables interprocedural (i.e., link-time) optimizations. Other code parameters are the alignment of buffers to transparent huge pages of 2 MiB and the loop unrolling factor set to 8. For GPU devices, we make use of the latest cuBLAS 10.2 library officially supported by the architecture vendor.